2x-7+6x=57 x=
-54=-7x+x+6 x=
x+2+7x=18 x=
2x+9=x+14 x=
10+2x=-4x+46 x=
2(1x+8)=4 x=
-4x-7=-4+x x=
5(2+6x)=10 x=
3(7-6x)=-123 x=

Answer by College Student(505)

Steps:
a. combine like terms
b. solve for x
.
2x-7+6x=57
8x=57+7
8x=64
x=64/8
x=8
.
-54=-7x+x+6
-54=-6x+6
-54+6=-6x
-48=-6x
-48/-6=x
8=x ...same as... x=8
.
x+2+7x=18
8x+2=18
8x=18-2
8x=16
x=16/8
x=2
.
2x+9=x+14
2x-x=14-9
x=5
.
10+2x=-4x+46
2x+4x=46-10
6x=36
x=36/6
x=6
.
In the next problem we need to distribute (multiply) 2 by each term inside the parenthesis, then combine like terms.
2(1x+8)=4
2*1x)+(2*8)=4
2x+16=4
2x=4-16
2x=-12
x=-12/2
x=-6
.
-4x-7=-4+x
-4x-x=-4+7
-5x=3
x=3/-5 ...same as to say... x=-3/5
.
Remember we need to distribute (multiply) 5 by each term inside the parenthesis.
5(2+6x)=10
(5*2)+(5*6x)=10
10+30x=10
30x=10-10
30x=0
x=0/30 ...zero divided by any number will always equal zero.
x=0
.
3(7-6x)=-123
(3*7)-(3*6x)=-123
21-18x=-123
18x=-123+21
18x=-102
x=-102/18 ... this fraction (or division - same thing) does not reduce to a whole number. In other words, if we take a calculator to divide -102/18, we get -5.66. We usually do not express these answers with decimals unless your teacher tells you it's ok.
.
Now, this fraction can be simplify to a smaller fraction and still mean the same thing. Since both numerator and denominator are even numbers, I would start by divding them by 2 and see what we get.
-102/2=-51
18/2=9
So we can say: x=-102/18 ...or... x=-51/9
.
Done!